OpenStudy (anonymous):

sin(theta)csc(theta)-sin(90-theta)sec(theta)=0

OpenStudy (anonymous):

\[\sin\theta\csc\theta-\sin(90-\theta)\sec\theta=0\] Use the angle difference identity for sine: \[\sin(x-y)=\sin x\cos y-\sin y\cos x\] Also, recall that \(\displaystyle\frac{1}{\sin\theta}=\csc\theta.\)